# Physics-Informed Teleconnection-Aware Transformer for Global Subseasonal-to-Seasonal Forecasting

### v2

@@ A.5.3 Generalization Bounds
The physics-inspired constraints effectively reduce the hypothesis space complexity for each frequency band, lowering $\color[rgb]{0.171875,0.2421875,0.3125}\mathcal{C}(h_{\ell})$ compared to unconstrained neural networks. The multi-scale decomposition enables specialized processing that further reduces per-band complexity. These reductions, combined with standard uniform convergence arguments, yield the stated bound. ∎
@@ B.1 Evaluation Metric
Following existing works liucirt ; nathaniel2024chaosbench ; rasp2024weatherbench , we adopt three metrics to evaluate the model performance: Root Mean Squared Error (RMSE) with a weighting scheme at each latitude $\color[rgb]{0.171875,0.2421875,0.3125}\theta_{i}$, Anomaly Correlation Coefficient (ACC), and Spectral Divergence (SpecDiv).
@@
$\text{RMSE}=\sqrt{\frac{1}{|\theta||\phi|}\sum_{i=1}^{|\theta|}\sum_{j=1}^{|\phi|}w(\theta_{i})(\hat{\mathbf{Y}}_{i,j}-\mathbf{Y}_{i,j})^{2}},\text{ACC}=\frac{\sum(\theta_{i})[A_{\hat{Y}}\cdot A_{Y}]}{\sqrt{\sum w(\theta_{i})A_{\hat{Y}}^{2}\sum w(\theta_{i})A_{Y}^{2}}},$ (23)
where $\color[rgb]{0.171875,0.2421875,0.3125}w(\theta_{i})=\frac{cos(\theta_{i})}{\frac{1}{|\theta|}\sum_{a=1}^{|\theta|}cos(\theta_{a})}$, $\color[rgb]{0.171875,0.2421875,0.3125}\theta_{i}\in\theta=[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\color[rgb]{0.171875,0.2421875,0.3125}\phi_{i}\in\phi=[-\pi,\pi]$ denotes the set of all latitudes and longitude, respectively. The predicted and observed anomalies at each grid are denoted by $\color[rgb]{0.171875,0.2421875,0.3125}A_{\hat{\mathbf{Y}}_{i,j}}=\hat{\mathbf{Y}}_{i,j}-C$ and $\color[rgb]{0.171875,0.2421875,0.3125}A_{\mathbf{Y}_{i,j}}=\mathbf{Y}_{i,j}-C$ where $\color[rgb]{0.171875,0.2421875,0.3125}C$ is observational climatology.
@@
For the physics-based metric, we adopt the SpecDiv as proposed in nathaniel2024chaosbench . This metric measures the deviation between the power spectra of prediction and target in the spectral domain. It is particularly useful for evaluating the preservation of high-frequency components in forecasts, which tend to become blurry due to power divergence in the spectral domain. The SpecDiv is defined as:
$\text{SpecDiv}=\sum_{k}S^{\prime}(k)\cdot\log(\frac{S^{\prime}(k)}{\hat{S}^{\prime}(k)})$ (24)
where $\color[rgb]{0.171875,0.2421875,0.3125}S^{\prime}(k)$ and $\color[rgb]{0.171875,0.2421875,0.3125}\hat{S}^{\prime}(k)$ represent the normalized power spectra of target and prediction respectively. This metric follows the principles of Kullback-Leibler divergence to quantify the spectral differences between predictions and ground truth.
@@
Let $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{Y}$ and $\color[rgb]{0.171875,0.2421875,0.3125}\hat{\mathbf{Y}}$ represent the ground-truth and predicted images respectively. MS-SSIM incorporates three comparison measures defined as:
$l(\mathbf{Y},\hat{\mathbf{Y}})=\frac{2\mu_{\mathbf{Y}}\mu_{\hat{\mathbf{Y}}}+C_{1}}{\mu_{\mathbf{Y}}^{2}+\mu_{\hat{\mathbf{Y}}}^{2}+C_{1}},c(\mathbf{Y},\hat{\mathbf{Y}})=\frac{2\sigma_{\mathbf{Y}}\sigma_{\hat{\mathbf{Y}}}+C_{2}}{\sigma_{\mathbf{Y}}^{2}+\sigma_{\hat{\mathbf{Y}}}^{2}+C_{2}},s(\mathbf{Y},\hat{\mathbf{Y}})=\frac{\sigma_{\mathbf{Y}\hat{\mathbf{Y}}}+C_{3}}{\sigma_{\mathbf{Y}}\sigma_{\hat{\mathbf{Y}}}+C_{3}}$ (25)
where $\color[rgb]{0.171875,0.2421875,0.3125}\mu_{\mathbf{Y}}$, $\color[rgb]{0.171875,0.2421875,0.3125}\sigma_{\mathbf{Y}}^{2}$ and $\color[rgb]{0.171875,0.2421875,0.3125}\sigma_{\mathbf{Y}\hat{\mathbf{Y}}}$ are the mean of $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{Y}$, the variance of $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{Y}$, and the covariance between $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{Y}$ and $\color[rgb]{0.171875,0.2421875,0.3125}\hat{\mathbf{Y}}$, respectively. The constants $\color[rgb]{0.171875,0.2421875,0.3125}C_{1}=(K_{1}L)^{2}$, $\color[rgb]{0.171875,0.2421875,0.3125}C_{2}=(K_{2}L)^{2}$, and $\color[rgb]{0.171875,0.2421875,0.3125}C_{3}=C_{2}/2$ are included to avoid instability when the denominators are close to zero, where $\color[rgb]{0.171875,0.2421875,0.3125}L=255$ is the dynamic range of the grayscale images, and $\color[rgb]{0.171875,0.2421875,0.3125}K_{1}\ll 1$ and $\color[rgb]{0.171875,0.2421875,0.3125}K_{2}\ll 1$ are two small constants. The MS-SSIM metric is computed by analyzing these components across multiple scales through successive low-pass filtering and downsampling. At each scale $\color[rgb]{0.171875,0.2421875,0.3125}j$, the contrast comparison $\color[rgb]{0.171875,0.2421875,0.3125}c_{j}(\mathbf{Y},\hat{\mathbf{Y}})$ and structure comparison $\color[rgb]{0.171875,0.2421875,0.3125}s_{j}(\mathbf{Y},\hat{\mathbf{Y}})$ are calculated, while luminance comparison $\color[rgb]{0.171875,0.2421875,0.3125}l_{M}(\mathbf{Y},\hat{\mathbf{Y}})$ is only considered at the coarsest scale $\color[rgb]{0.171875,0.2421875,0.3125}M$. The final MS-SSIM metric is defined as:
@@
where $\color[rgb]{0.171875,0.2421875,0.3125}\alpha_{M}$, $\color[rgb]{0.171875,0.2421875,0.3125}\beta_{j}$, and $\color[rgb]{0.171875,0.2421875,0.3125}\gamma_{j}$ are weighting parameters. We follow standard practice by setting $\color[rgb]{0.171875,0.2421875,0.3125}M=5$ and using the parameter values established in previous literaturenathaniel2024chaosbench .
@@
$\text{SpecRes}=\sqrt{\mathbb{E}_{k}[(\hat{S}^{\prime}(k)-S^{\prime}(k))^{2}]}$ (27)
where $\color[rgb]{0.171875,0.2421875,0.3125}\hat{S}^{\prime}(k)$ and $\color[rgb]{0.171875,0.2421875,0.3125}S^{\prime}(k)$ are the normalized power spectra of prediction and target respectively, and $\color[rgb]{0.171875,0.2421875,0.3125}\mathbb{E}_{k}$ denotes expectation over the high-frequency wavenumbers $\color[rgb]{0.171875,0.2421875,0.3125}\mathbf{K}_{q}$. Similar to SpecDiv, this metric will be zero if the power spectra are identical, but it emphasizes the magnitude of spectral differences rather than their distributional divergence.
B.2 Baselines
@@
B.3 Additional Overall Performance
@@
Table 5 extends our comparison against operational forecasting systems across all five metrics (RMSE, ACC, MS-SSIM, SpecDiv, and SpecRes). The results demonstrate TelePiT’s significant performance advantages over current operational systems across all metrics and forecast horizons, with improvements ranging from 20-40% depending on the variable and metric.
@@
B.4 Additional Ablation Study
@@
B.5 Additional Empirical Analysis
@@ B.6 Robustness Study
B.7 Parameter Sensitivity Analysis
Figure 20 shows the sensitivity analysis of teleconnection coefficient $\color[rgb]{0.171875,0.2421875,0.3125}\lambda$, which controls the balance between local dynamics and global teleconnection patterns in the attention mechanism. The results reveal a consistent optimal value of $\color[rgb]{0.171875,0.2421875,0.3125}\lambda=0.2$ across most variables and metrics, where RMSE reaches minimum and ACC achieves maximum values. Temperature variables show sharp optima, while wind components exhibit the most pronounced sensitivity, reflecting the complex interplay between boundary layer processes and large-scale pressure gradients. The consistency of $\color[rgb]{0.171875,0.2421875,0.3125}\lambda=0.2$ across different variables validates our architectural design and simplifies model deployment.
Table 4: Performance comparison of TelePiT against data-driven models: MS-SSIM (higher values indicate better structural similarity) and SpecRes (lower values indicate better spectral accuracy). Results are shown for both Weeks 3-4 and Weeks 5-6 forecasts, demonstrating TelePiT’s consistent performance advantages across all key variables, metrics, and forecast lead times.
@@ B.7 Parameter Sensitivity Analysis
Table 5: Performance comparison of TelePiT against operational forecasting systems. Results are shown for both Weeks 3-4 and 5-6 forecasts, demonstrating TelePiT’s significant performance advantages over current operational systems across all metrics.
@@
RMSE ($\downarrow$) UKMO 64.2965 43.1836 2.1440 2.5763 63.0383 45.6263 2.2743 2.3924
NCEP 68.4408 44.7869 2.2250 2.7422 71.3320 46.8340 2.3433 2.8331
ECMWF 63.6737 42.3491 2.0667 2.3910 68.1156 45.3303 2.1772 2.5110
CMA 69.8326 47.0562 2.5975 2.9181 71.3996 48.5830 2.6388 2.9334
TelePiT 48.6714 31.0823 1.6849 1.8739 47.7618 30.4056 1.6985 1.8949
ACC ($\uparrow$) UKMO 0.9726 0.9294 0.9804 0.9768 0.9695 0.9225 0.9781 0.9767
NCEP 0.9694 0.9206 0.9790 0.9774 0.9669 0.9133 0.9766 0.9760
ECMWF 0.9724 0.9289 0.9815 0.9828 0.9677 0.9157 0.9794 0.9807
CMA 0.9679 0.9188 0.9746 0.9744 0.9660 0.9121 0.9739 0.9738
TelePiT 0.9848 0.9626 0.9882 0.9900 0.9852 0.9643 0.9881 0.9897
MS-SSIM ($\uparrow$) UKMO 0.8682 0.7959 0.8848 0.8004 0.8572 0.7767 0.8747 0.8570
@@
SpecDiv ($\downarrow$) UKMO 0.1173 0.1142 0.2408 0.0989 0.1330 0.1267 0.2905 0.1182
NCEP 0.5567 0.4875 4.3344 0.6459 0.5598 0.4826 4.2610 0.6135
ECMWF 0.3976 0.0599 0.1229 0.2375 0.3093 0.0687 0.0575 0.1187
CMA 0.1612 0.1671 0.2725 0.0318 0.1589 0.1544 0.1277 0.0324
TelePiT 0.0180 0.0175 0.0150 0.0045 0.0161 0.0168 0.0184 0.0042
SpecRec ($\downarrow$) UKMO 0.0386 0.0377 0.0470 0.0310 0.0681 0.0403 0.0523 0.0302
@@
Figure 6: ACC comparison of variable $\color[rgb]{0.171875,0.2421875,0.3125}z$, $\color[rgb]{0.171875,0.2421875,0.3125}t$, $\color[rgb]{0.171875,0.2421875,0.3125}u$, and $\color[rgb]{0.171875,0.2421875,0.3125}v$ on all different pressure levels. The color scale indicates normalized ACC values (higher is better), with darker blue representing superior performance. TelePiT consistently shows higher ACC values across all variables and pressure levels compared to both data-driven models and operational forecasting systems. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Figure 7: MS-SSIM comparison of variable $\color[rgb]{0.171875,0.2421875,0.3125}z$, $\color[rgb]{0.171875,0.2421875,0.3125}t$, $\color[rgb]{0.171875,0.2421875,0.3125}u$, and $\color[rgb]{0.171875,0.2421875,0.3125}v$ on all different pressure levels. The color gradient indicates normalized MS-SSIM values (higher is better), showing TelePiT’s superior ability to preserve structural features in forecasts compared to all baseline models, with particularly strong performance in temperature and wind variables at lower-to-mid tropospheric levels. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Figure 8: SpecDiv comparison of variable $\color[rgb]{0.171875,0.2421875,0.3125}z$, $\color[rgb]{0.171875,0.2421875,0.3125}t$, $\color[rgb]{0.171875,0.2421875,0.3125}u$, and $\color[rgb]{0.171875,0.2421875,0.3125}v$ on all different pressure levels. This metric quantifies differences in power spectrum distributions between predicted and ground truth fields. Lower values (lighter colors) indicate better spectral match. TelePiT demonstrates consistently lower SpecDiv values across variables and pressure levels, highlighting its ability to accurately capture atmospheric wave characteristics at various spatial scales. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Figure 9: SpecRes comparison of variable $\color[rgb]{0.171875,0.2421875,0.3125}z$, $\color[rgb]{0.171875,0.2421875,0.3125}t$, $\color[rgb]{0.171875,0.2421875,0.3125}u$, and $\color[rgb]{0.171875,0.2421875,0.3125}v$ on all different pressure levels. This metric measures spectral energy distribution accuracy. Lower values (lighter colors) indicate better performance. TelePiT exhibits substantially lower SpecRes values compared to all competing models, with particular advantages in temperature and wind fields, demonstrating its enhanced ability to represent both large-scale circulation patterns and finer-scale atmospheric features. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Table 6: Extended ablation study comparing TelePiT against variants with removed components (SHE: Spherical Harmonic Embedding, WD: Wavelet Decomposition, ODE: Physics-Informed ODE, TA: Teleconnection Attention) using MS-SSIM (higher is better), SpecDiv and SpecRec (lower is better) metrics across various atmospheric variables for both Weeks 3-4 and 5-6 forecast horizons.
@@
Figure 10: Global visualization of RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}t850$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 11: Global visualization of RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}t500$ for Weeks 3-4 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error).
@@
Figure 12: Global visualization of RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}t500$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 13: Global visualization of RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}z850$ for Weeks 3-4 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error).
@@
Figure 14: Global visualization of RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}z850$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 15: Global visualization of RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}z500$ for Weeks 3-4 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error).
@@
Figure 16: Global visualization of RMSE distribution of $\color[rgb]{0.171875,0.2421875,0.3125}z500$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 17: Daily forecasting performance (RMSE) of $\color[rgb]{0.171875,0.2421875,0.3125}t500$ in 2018. For improved visualization clarity, a Gaussian smoothing (with $\color[rgb]{0.171875,0.2421875,0.3125}\sigma=1.5$) was applied to the daily values for models exhibiting high variability (Pangu, GraphCast, FourCastNetV2, CMA, ECMWF, NCEP, and UKMO). Note that ECMWF predictions appear as discrete points because their operational system generates forecasts only every 2-3 days, unlike the other models which provide daily predictions.
Figure 18: Daily forecasting performance (RMSE) of $\color[rgb]{0.171875,0.2421875,0.3125}z850$ in 2018.
Figure 19: Daily forecasting performance (RMSE) of $\color[rgb]{0.171875,0.2421875,0.3125}z500$ in 2018.
@@
Table 7: Robustness study on 2019 data (out-of-sample) comparing TelePiT against baselines across all metrics, demonstrating consistent performance advantages and strong generalization capabilities. Operational systems (NCEP, UKMO, CMA, ECMWF) results are obtained from ChaosBench and do not include $t2m$, $u10$, $v10$ predictions. GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues.
@@
Figure 20: Parameter sensitivity analysis of teleconnection coefficient $\color[rgb]{0.171875,0.2421875,0.3125}\lambda$ across different key variables and all evaluation metrics on Weeks 3-4. The parameter $\color[rgb]{0.171875,0.2421875,0.3125}\lambda$ controls the influence of teleconnection patterns in the attention mechanism (Equation 8). Most variables achieve optimal performance at $\color[rgb]{0.171875,0.2421875,0.3125}\lambda$ = 0.2, indicating that moderate incorporation of atmospheric teleconnection patterns enhances model performance. RMSE values generally reach minimum at $\color[rgb]{0.171875,0.2421875,0.3125}\lambda$ = 0.2, while ACC values peak at the same point, suggesting an optimal balance between local atmospheric dynamics and large-scale teleconnection patterns for S2S forecasting accuracy.

### v3

@@ A.5.3. Generalization Bounds
The physics-inspired constraints effectively reduce the hypothesis space complexity for each frequency band, lowering $\mathcal{C}(h_{\ell})$ compared to unconstrained neural networks. The multi-scale decomposition enables specialized processing that further reduces per-band complexity. These reductions, combined with standard uniform convergence arguments, yield the stated bound. ∎
@@ B.1. Evaluation Metric
Following existing works (Liu et al., 2025; Nathaniel et al., 2024; Rasp et al., 2024), we adopt three metrics to evaluate the model performance: Root Mean Squared Error (RMSE) with a weighting scheme at each latitude $\theta_{i}$, Anomaly Correlation Coefficient (ACC), and Spectral Divergence (SpecDiv).
@@
(35) $\text{RMSE}=\sqrt{\frac{1}{|\theta||\phi|}\sum_{i=1}^{|\theta|}\sum_{j=1}^{|\phi|}w(\theta_{i})(\hat{\mathbf{Y}}_{i,j}-\mathbf{Y}_{i,j})^{2}},$
(36) $\text{ACC}=\frac{\sum(\theta_{i})[A_{\hat{Y}}\cdot A_{Y}]}{\sqrt{\sum w(\theta_{i})A_{\hat{Y}}^{2}\sum w(\theta_{i})A_{Y}^{2}}},$
where $w(\theta_{i})=\frac{cos(\theta_{i})}{\frac{1}{|\theta|}\sum_{a=1}^{|\theta|}cos(\theta_{a})}$, $\theta_{i}\in\theta=[-\frac{\pi}{2},\frac{\pi}{2}]$ and $\phi_{i}\in\phi=[-\pi,\pi]$ denotes the set of all latitudes and longitude, respectively. The predicted and observed anomalies at each grid are denoted by $A_{\hat{\mathbf{Y}}_{i,j}}=\hat{\mathbf{Y}}_{i,j}-C$ and $A_{\mathbf{Y}_{i,j}}=\mathbf{Y}_{i,j}-C$ where $C$ is observational climatology.
@@
For the physics-based metric, we adopt the SpecDiv as proposed in (Nathaniel et al., 2024). This metric measures the deviation between the power spectra of prediction and target in the spectral domain. It is particularly useful for evaluating the preservation of high-frequency components in forecasts, which tend to become blurry due to power divergence in the spectral domain. The SpecDiv is defined as:
(37) $\text{SpecDiv}=\sum_{k}S^{\prime}(k)\cdot\log(\frac{S^{\prime}(k)}{\hat{S}^{\prime}(k)})$
where $S^{\prime}(k)$ and $\hat{S}^{\prime}(k)$ represent the normalized power spectra of target and prediction respectively. This metric follows the principles of Kullback-Leibler divergence to quantify the spectral differences between predictions and ground truth.
@@
Let $\mathbf{Y}$ and $\hat{\mathbf{Y}}$ represent the ground-truth and predicted images respectively. MS-SSIM incorporates three comparison measures defined as:
(38) $l(\mathbf{Y},\hat{\mathbf{Y}})=\frac{2\mu_{\mathbf{Y}}\mu_{\hat{\mathbf{Y}}}+C_{1}}{\mu_{\mathbf{Y}}^{2}+\mu_{\hat{\mathbf{Y}}}^{2}+C_{1}},c(\mathbf{Y},\hat{\mathbf{Y}})=\frac{2\sigma_{\mathbf{Y}}\sigma_{\hat{\mathbf{Y}}}+C_{2}}{\sigma_{\mathbf{Y}}^{2}+\sigma_{\hat{\mathbf{Y}}}^{2}+C_{2}},s(\mathbf{Y},\hat{\mathbf{Y}})=\frac{\sigma_{\mathbf{Y}\hat{\mathbf{Y}}}+C_{3}}{\sigma_{\mathbf{Y}}\sigma_{\hat{\mathbf{Y}}}+C_{3}}$
@@
(39) $\text{MS-SSIM}=[l_{M}(\mathbf{Y},\hat{\mathbf{Y}})]^{\alpha_{M}}\cdot\prod_{j=1}^{M}[c_{j}(\mathbf{Y},\hat{\mathbf{Y}})]^{\beta_{j}}[s_{j}(\mathbf{Y},\hat{\mathbf{Y}})]^{\gamma_{j}}$
where $\alpha_{M}$, $\beta_{j}$, and $\gamma_{j}$ are weighting parameters. We follow standard practice by setting $M=5$ and using the parameter values established in previous literature(Nathaniel et al., 2024).
@@
(40) $\text{SpecRes}=\sqrt{\mathbb{E}_{k}[(\hat{S}^{\prime}(k)-S^{\prime}(k))^{2}]}$
where $\hat{S}^{\prime}(k)$ and $S^{\prime}(k)$ are the normalized power spectra of prediction and target respectively, and $\mathbb{E}_{k}$ denotes expectation over the high-frequency wavenumbers $\mathbf{K}_{q}$. Similar to SpecDiv, this metric will be zero if the power spectra are identical, but it emphasizes the magnitude of spectral differences rather than their distributional divergence.
B.2. Baselines
@@
B.3. Model Complexity
We analyze the computational complexity of TelePiT compared to state-of-the-art baselines in Table 5. While TelePiT requires more resources than the lightweight CirT (14.5G vs 2.2G FLOPs, 37M vs 16M parameters), this modest increase delivers substantial accuracy improvements, reducing forecast errors by up to 20% across critical variables. More importantly, TelePiT remains dramatically more efficient than current leading models, requiring 7,500× fewer FLOPs than GraphCast (110T) and 11,500× fewer than PanguWeather (168T). With a practical model size of 141.64 MB and real-time inference capability (32.25 ms, 31 samples/second), TelePiT can be deployed on standard GPUs or even high-end edge devices—a critical advantage for operational weather services worldwide. This optimal balance between accuracy and efficiency stems from our physics-informed architecture that replaces brute-force spatial computation with intelligent design: spherical harmonic embedding reduces complexity from $O(H\times W)$ to $O(H)$, while teleconnection-aware attention leverages sparse global patterns instead of dense pixel-wise interactions. For S2S forecasting where both accuracy and computational feasibility determine real-world impact, TelePiT demonstrates that incorporating domain knowledge can achieve state-of-the-art performance at a fraction of the computational cost, making advanced climate prediction accessible even in resource-constrained settings.
Table 5. Computation complexity and model size comparison.
Model GraphCast PanguWeather CirT TelePiT
FLOPs 110T 168T 2.2G 14.5G
Params 37M 256M 16M 37M
B.4. Additional Overall Performance
@@
Table 7 extends our comparison against operational forecasting systems across all five metrics (MS-SSIM and SpecRes). The results demonstrate TelePiT’s significant performance advantages over current operational systems across all metrics and forecast horizons, with improvements ranging from 20-40% depending on the variable and metric.
@@
B.5. Additional Ablation Study
@@
B.6. Additional Empirical Analysis
@@ B.7. Robustness Study
B.8. Parameter Sensitivity Analysis
Figure 20 shows the sensitivity analysis of teleconnection coefficient $\lambda$, which controls the balance between local dynamics and global teleconnection patterns in the attention mechanism. The results reveal a consistent optimal value of $\lambda=0.2$ across most variables and metrics, where RMSE reaches minimum and ACC achieves maximum values. Temperature variables show sharp optima, while wind components exhibit the most pronounced sensitivity, reflecting the complex interplay between boundary layer processes and large-scale pressure gradients. The consistency of $\lambda=0.2$ across different variables validates our architectural design and simplifies model deployment.
Table 6. Performance comparison of TelePiT against data-driven models: MS-SSIM (higher values indicate better structural similarity) and SpecRes (lower values indicate better spectral accuracy). Results are shown for both Weeks 3-4 and Weeks 5-6 forecasts, demonstrating TelePiT’s consistent performance advantages across all key variables, metrics, and forecast lead times.
@@ B.8. Parameter Sensitivity Analysis
Table 7. Performance comparison of TelePiT against operational forecasting systems. Results are shown for both Weeks 3-4 and 5-6 forecasts, demonstrating TelePiT’s significant performance advantages over current operational systems across all metrics.
@@
MS-SSIM ($\uparrow$)
UKMO 0.8682 0.7959 0.8848 0.8004 0.8572 0.7767 0.8747 0.8570
@@
SpecRec ($\downarrow$)
UKMO 0.0386 0.0377 0.0470 0.0310 0.0681 0.0403 0.0523 0.0302
@@
Figure 6. ACC comparison of variable $z$, $t$, $u$, and $v$ on all different pressure levels. The color scale indicates normalized ACC values (higher is better), with darker blue representing superior performance. TelePiT consistently shows higher ACC values across all variables and pressure levels compared to both data-driven models and operational forecasting systems. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Figure 7. MS-SSIM comparison of variable $z$, $t$, $u$, and $v$ on all different pressure levels. The color gradient indicates normalized MS-SSIM values (higher is better), showing TelePiT’s superior ability to preserve structural features in forecasts compared to all baseline models, with particularly strong performance in temperature and wind variables at lower-to-mid tropospheric levels. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Figure 8. SpecDiv comparison of variable $z$, $t$, $u$, and $v$ on all different pressure levels. This metric quantifies differences in power spectrum distributions between predicted and ground truth fields. Lower values (lighter colors) indicate better spectral match. TelePiT demonstrates consistently lower SpecDiv values across variables and pressure levels, highlighting its ability to accurately capture atmospheric wave characteristics at various spatial scales. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Figure 9. SpecRes comparison of variable $z$, $t$, $u$, and $v$ on all different pressure levels. This metric measures spectral energy distribution accuracy. Lower values (lighter colors) indicate better performance. TelePiT exhibits substantially lower SpecRes values compared to all competing models, with particular advantages in temperature and wind fields, demonstrating its enhanced ability to represent both large-scale circulation patterns and finer-scale atmospheric features. GraphCast (GC) is unavailable at Weeks 5-6 due to the out of memory.
Table 8. Extended ablation study comparing TelePiT against variants with removed components (SHE: Spherical Harmonic Embedding, WD: Wavelet Decomposition, ODE: Physics-Informed ODE, TA: Teleconnection Attention) using MS-SSIM (higher is better), SpecDiv and SpecRec (lower is better) metrics across various atmospheric variables for both Weeks 3-4 and 5-6 forecast horizons.
@@
Figure 10. Global visualization of RMSE distribution of $t850$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 11. Global visualization of RMSE distribution of $t500$ for Weeks 3-4 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error).
@@
Figure 12. Global visualization of RMSE distribution of $t500$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 13. Global visualization of RMSE distribution of $z850$ for Weeks 3-4 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error).
@@
Figure 14. Global visualization of RMSE distribution of $z850$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 15. Global visualization of RMSE distribution of $z500$ for Weeks 3-4 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error).
@@
Figure 16. Global visualization of RMSE distribution of $z500$ for Weeks 5-6 forecasts across different models. The color scale represents RMSE values, with darker blue indicating better performance (lower error) and darker red indicating worse performance (higher error). Note that GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues when attempting extended-range forecasts.
@@
Figure 17. Daily forecasting performance (RMSE) of $t500$ in 2018. For improved visualization clarity, a Gaussian smoothing (with $\sigma=1.5$) was applied to the daily values for models exhibiting high variability (Pangu, GraphCast, FourCastNetV2, CMA, ECMWF, NCEP, and UKMO). Note that ECMWF predictions appear as discrete points because their operational system generates forecasts only every 2-3 days, unlike the other models which provide daily predictions.
Figure 18. Daily forecasting performance (RMSE) of $z850$ in 2018.
Figure 19. Daily forecasting performance (RMSE) of $z500$ in 2018.
@@
Table 9. Robustness study on 2019 data (out-of-sample) comparing TelePiT against baselines across all metrics, demonstrating consistent performance advantages and strong generalization capabilities. Operational systems (NCEP, UKMO, CMA, ECMWF) results are obtained from ChaosBench and do not include $t2m$, $u10$, $v10$ predictions. GraphCast results are unavailable at Weeks 5-6 due to out-of-memory issues.
@@
Figure 20. Parameter sensitivity analysis of teleconnection coefficient $\lambda$ across different key variables and all evaluation metrics on Weeks 3-4. The parameter $\lambda$ controls the influence of teleconnection patterns in the attention mechanism (Equation 8). Most variables achieve optimal performance at $\lambda$ = 0.2, indicating that moderate incorporation of atmospheric teleconnection patterns enhances model performance. RMSE values generally reach minimum at $\lambda$ = 0.2, while ACC values peak at the same point, suggesting an optimal balance between local atmospheric dynamics and large-scale teleconnection patterns for S2S forecasting accuracy.